# MEBench: Benchmarking Large Language Models for Cross-Document Multi-Entity Question Answering

### v1

@@ 3.2.2. Question and answer Generation
Question Generation
The foundational input for the QA generation pipeline is the table generated in last step. The generation of questions follows a ”template-driven, entity-attribute coupling” paradigm, implemented through LLM (GPT-4), as illustrated in Figure 2-c1. Predefined syntactic and semantic templates govern the grammatical structure and intent of questions. These templates are shown in Appendix Table 6. The LLM instantiates templates with entity-attribute pairs, ensuring syntactic diversity while adhering to logical constraints. Generated questions undergo validation via: Syntactic Checks, Ensuring grammatical correctness; Semantic Grounding, verifying that the question is answerable using the table’s data. Ambiguity Reduction, pruning underspecified questions (e.g., ”Describe the economy” revised to ”Describe the GDP growth rate of Brazil in 2023”).
Answer Generation
Answers are derived programmatically through automated SQL query execution, ensuring reproducibility and alignment with the table’s ground-truth data. The synthesized SQL is executed against the table, yielding direct answers or sub-tables (Intermediate results requiring post-processing), as illustrated in Figure 2-c3. Answers are standardized to ensure consistency: Numeric results are rounded to significant figures; Categorical answers are converted to canonical forms (e.g., ”USA” to ”United States”).
Integration and Validation
The final output consists of a dataset where each question is matched with its corresponding answer derived from SQL queries, thereby meeting the specified requirements:
  • Traceability: Each answer is directly linked to the table via its SQL provenance.
  • Scalability: The template-driven approach supports rapid generation of diverse QA pairs.
  • Rigor: Execution-based answering eliminates hallucination risks inherent in generative-only methods.
3.2.3. Quality Control
Question Templates.
The use of templates ensures that every question is crafted with a clear structure, making it easier for respondents to understand and answer them accurately. For relationship and complex statistic questions we turn the questions in a closed-ended style, as they require a specific response of either ”yes” or ”no”, which make the answer in a standardized format. The examples of Question Templates is in the Appendix 6.
Question Refinement.
Manual review.
We assess the questions for accuracy, ensuring they are factually correct and relevant to the our purpose. Manual reviews can also provide insights into whether the questions are likely to effectively elicit the intended information from answers, thereby contributing to the reliability and validity of the benchmark.
3.3. Data Statistics
The benchmark comprises 4,780 methodically structured questions partitioned into two subsets: a training set (3,406 questions) for model fine-tuning and a test set (1,374 questions) for rigorous evaluation. Based on entity count, the data is divided into three groups: “low” (0-10), “Medium” (11-100), and “high” (¿100), containing 683, 1366, and 2731 entries, respectively. Table 2 details comprehensive statistics of the benchmark. We also analyze the proportion of questions rejected during manual review and find that 21% of the questions are rejected for unqualified question.
4. Experiment
Table 3. Experimental results for MEBench.
@@ 4. Experiment
Table 4. Quality of Large Language Models (LLMs) in EA-F1.
Models $EA-F1$
@@
Figure 3. The Experimental results for eight subtasks of each model.
4.1. Experiment Setup
4.1.1. Models
For open-source LLMs, we conduct experiments using the representative Meta-Llama-3-8B-Instruct (Meta Llama3, 2024) and apply QLoRA (Dettmers et al., 2023) to fine-tune it with the training set of MEBench. For proprietary LLMs, we select the widely recognized GPT models, including GPT-3.5-turbo (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023). Additionally, we incorporate RAG across all vanilla baseline models for comparative analysis and evaluation of the model’s capacity to integrate and leverage external data sources.
4.1.2. Structured RAG
We have also incorporated a structured RAG module into the LLMs to explore whether RAG can enhance the model’s performance on MEBench. For the Embedding choice, we employ the OpenAI Embedding model (Tex, [n. d.]), and the chunk size is 1024. For each document, we retrieve the top-5 most related chunks and concatenate them in their original order to form the context input for the model.
4.1.3. Evaluation Metrics
We adopt Accuracy ($Acc$) as the primary metric to assess the performance of LLMs on MEBench tasks. For the subcategories of Variance Analysis, Correlation Analysis, and Distribution Compliance within the Statistics tasks shown in Table 1, we focus solely on prompting LLMs to identify relevant columns and applicable methods, evaluating the accuracy of their selections instead of the computational results, as LLMs’ abilities in precise calculations are not the central focus of this study. In addition, we evaluate performance of information extraction using Entity-Attributed F1 (EA-F1). This is an F1 score applied to the predicted vs. gold sets of the (entity, atrribution, value). All three elements in the tuple must exactly match the tuple in the ground truth to be marked correct.
4.2. Results and Analysis
@@
4.2.1. Main result
GPT-4 + RAG achieved superior accuracy (59.3%), outperforming the second-ranked model (FT Llama-3-Instruct: 55.6% ) by a statistically significant margin. Notably, GPT-4 + RAG excelled in relational (68.7%) and comparative (76.3%) queries, likely due to its superior contextual understanding. However, all models exhibited markedly lower accuracy in statistical queries (GPT-4 + RAG: 41.0%), suggesting inherent challenges in numerical reasoning. In our evaluation, we focused on analyzing the capability of LLMs to extract related data. This assessment aimed to understand how well these sophisticated models can organize and present data for the question. The result is shown in Table 4. These results underscore the critical role of information extraction architectures in mitigating hallucinations and grounding outputs in factual data. Introducing RAG significantly improves overall performance, particularly in comparison tasks, while fine-tuning LLaMA-3-Instruct alone does not yield substantial gains without RAG. On MEQA, open-source models like LLaMA-3-Instruct, even with RAG, can’t match proprietary models like GPT-4, which achieves a 59.3% accuracy compared to LLaMA-3-Instruct’s 31.6%.
4.2.2. Fine-grained Performance on Sub-tasks.
Figure 3 shows that vanilla LLMs perform well in correlation analysis and descriptive relationship sub-tasks, while RAG significantly improves intercomparison and superlative tasks. However, neither fine-tuning nor RAG overcomes challenges in variance analysis and aggregation tasks, while GPT-4 + RAG achieves accuracy of 15.3% and 32.1%.
4.2.3. Entity density Analysis.
As we can see from Table 3, our experiments underscore the impact of entity density on model performance in MEQA tasks. This phenomenon arises because higher entity densities amplify two critical challenges inherent to MEQA systems: (1) Semantic ambiguity due to overlapping relational predicates among entities (e.g., distinguishing ”Paris [person]” vs. ”Paris [location]” within narrow contexts), and (2) computational overhead in attention-based architectures attempting parallel reasoning over entangled entity-attribution pairs (e.g. transformer self-attention weights saturate under dense cross-entity dependencies).
  • Low Entity Density: Models generally performed well in low-density scenarios. The simplicity of context allowed for accurate entity recognition and minimal ambiguity in resolving references.
@@ 4.2.1. Main result
5. Conclusion
In this study, we have comprehensively addressed the significant challenges that multi-entity question answering (MEQA) poses to large language models (LLMs) and retrieval-augmented generation (RAG) systems. The limitations of existing methods in handling cross-document aggregation, especially when dealing with entity - dense questions, have been clearly identified and analyzed. We introduced MEBench, a groundbreaking multi-document, multi-entity benchmark. By systematically categorizing 4,780 questions into three primary categories and eight distinct types, MEBench offers broad coverage of real-world multi-entity reasoning scenarios. This categorization is a crucial step in providing a structured and comprehensive evaluation framework for LLMs. Our experiments on state-of-the-art LLMs such as GPT-4 and Llama-3, along with RAG pipelines, have shed light on the critical limitations of these advanced models. The fact that even these leading models achieve only 59% accuracy on MEBench underscores the magnitude of the challenges in MEQA. MEBench has effectively highlighted the systemic weaknesses in current LLM frameworks. These weaknesses serve as valuable insights for future research directions. For instance, the need for improved algorithms to retrieve and consolidate fragmented information from heterogeneous sources is evident. Additionally, there is a pressing need to develop more robust entity-aware QA architectures that can better handle the complexities of MEQA.
@@ Appendix A Appendix
A.1. Methodology for composite SPARQL Generation via Iterative Semantic Refinement
A.1.1. Initial Query Parsing Using GPT-4
Entity Misalignment: Incorrect Wikidata Q-ID assignments due to lexical ambiguity (e.g., ”Java” as programming language vs. Indonesian island)
A.1.2. Semantic Validation Layer
(a) Structured Knowledge Anchoring
The system interfaces with the Wikipedia API through programmatic endpoints that map surface forms to canonical entities via: {mdframed} def getwikidataid(term):
response = requests.get(
f”https:en.wikipedia.orgwapi.php?
action=query&format=json&prop=pageprops&titles=term”
return response.json()[”query”][”pages”].get(”pageprops”,
).get(”wikibaseitem”) (b) Neural-Semantic Disambiguation Module
A fine-tuned GPT-4 variant serves as our semantic analysis engine, performing three key operations: 1. Contextual disambiguation using entity linking algorithms enhanced by Wikifier-style mention detection 2. Property type validation against Wikidata’s ontology constraints (rdf:type, owl:ObjectProperty) 3. Temporal scope verification for time-sensitive queries requiring qualifiers like P585 [point in time]
A.1.3. Iterative Refinement Protocol
Executing candidate SPARQL on the Wikidata Query Service endpoint Analyzing result cardinality and type consistency 3 Applying constraint satisfaction heuristics: FILTER (?population ¿ 1e6 && ?country IN wd:Q30) # Example numerical/entity constraints Each iteration tightens precision metrics until meeting termination criteria defined by either:
$\frac{|ValidResults_{t}|}{|TotalResults_{t}|}\geq\theta_{precision}\quad(\theta=0.98\;\text{empirically})$
or maximum iteration thresholds.
A.1.4. Final Query Synthesis
A.2. Optimization
Model Selection.
Model selection is straightforward yet highly effective for optimization (Liu et al., 2024). Our system comprises multiple tasks, necessitating the selection of the most suitable model for different tasks. For basic tasks, more affordable and faster LLMs can suffice, while utilization of the most advanced LLMs is essential in more complex tasks to ensure optimal performance. Specifically, our system employs powerful yet resource-intensive GPT-4 for tasks such as semantic analysis or generation of table schemas and SQL queries. In contrast, for more basic information extraction, we utilize open-source Mistral-7B, thereby achieving a balance between cost efficiency and functional performance.
LLM Input/Output Control
SplitWise (Patel et al., 2023) shows that LLM inference time is generally proportional to the size of input and output tokens. Since GPT models decide the cost based on the input token, we try to minimize the input of large models. Meanwhile, we use the instructive prompt to reduce the size of the outputs generated by LLM without changing the quality of these outputs. The example of prompt is in Appendix A.2.1.
A.2.1. Prompt for Output Control
{mdframed}
A.3. Question Templates
By encoding each interrogative pattern into modular syntactic frames (e.g. ”Does [Entity X] exhibit [Attribute Y] when [Condition Z]?”), we achieve three critical objectives: (1) Elimination of lexical ambiguities through controlled vocabulary substitution ; (2) predictable semantic scaffolding that primes respondents toward domain-appropriate reasoning pathways; (3) standardized response elicitation critical for comparative analytics across heterogeneous samples.
A.4. Manual review
The manual review process constitutes a critical quality control mechanism in benchmark development, systematically implemented through three key phases: Content Validation Protocol, Psychometric Evaluation and Reliability Optimization. A tiered review system enhances inter-rater reliability: a) Primary coding by two independent reviewers b) Discrepancy resolution via consensus panels c) Final calibration against reference standards.
A.5. Tables
Table 5. Example Topics and Their Entities Attributions.
@@ A.5. Tables
Table 6. Template example for questions generated by the LLM (GPT-4).

### v2

@@ 3.2.2 Question and answer Generation
Question Generation.
The foundational input for the QA generation pipeline is the table generated in last step. The generation of questions follows a "template-driven, entity-attribute coupling" paradigm, implemented through LLM (GPT-4), as illustrated in Figure 2-c1. Predefined syntactic and semantic templates govern the grammatical structure and intent of questions. These templates are shown in Appendix Table 6. The LLM instantiates templates with entity-attribute pairs, ensuring syntactic diversity while adhering to logical constraints. Generated questions undergo validation via: Syntactic checks, ensuring grammatical correctness; Semantic grounding, verifying that the question is answerable using the table’s data; Ambiguity reduction, pruning underspecified questions (e.g., “Describe the economy” revised to “Describe the GDP growth rate of Brazil in 2023”).
Answer Generation.
Answers are derived programmatically through automated SQL query execution, ensuring reproducibility and alignment with the table’s ground-truth data. The synthesized SQL is executed against the table, yielding direct answers or sub-tables (Intermediate results requiring post-processing), as illustrated in Figure 2-c3. Answers are standardized to ensure consistency: Numeric results are rounded to significant figures; Categorical answers are converted to canonical forms (e.g., "USA" to "United States").
3.3 Data Statistics
The benchmark comprises 4,780 methodically structured questions partitioned into two subsets: a training set (3,406 questions) for model fine-tuning or train, and a test set (1,374 questions) for rigorous evaluation. Based on entity count, the data is divided into three groups: “low” (0-10), “Medium” (11-100), and “high” (>100), containing 683, 1366, and 2731 entries, respectively. Table 2 details comprehensive statistics of the benchmark. We also analyze the proportion of questions rejected during manual review and about 21% of the questions are failure to meet quality standards.
4 Experiment
Table 3: Experimental results for MEBench.
@@ 4 Experiment
Table 4: Quality of Large Language Models (LLMs) in EA-F1.
Models EAF1EA-F1
@@
Figure 3: The Experimental results for eight subtasks of each model.
4.1 Experiment Setup
Models. For open-source LLMs, we conduct experiments using the representative Meta-Llama-3-8B-Instruct (Meta Llama3, 2024) and apply QLoRA (Dettmers et al., 2023) to fine-tune it with the training set of MEBench. For proprietary LLMs, we select the widely recognized GPT models, including GPT-3.5-turbo (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023).
RAG. We implement a hierarchical retrieval framework that explicitly incorporates document organizational structures into the RAG pipeline to explore whether RAG can enhance the model’s performance on MEBench. For the Embedding choice, we employ the OpenAI Embedding model (OpenAI, ), and the chunk size is 1024. For each document, we retrieve the top-5 most related chunks and concatenate them in their original order to form the context input for the model.
Evaluation Metrics. We adopt Accuracy ($Acc$) as the primary metric to assess the performance of LLMs on MEBench tasks. For the subcategories of Variance Analysis, Correlation Analysis, and Distribution Compliance within the Statistics tasks, which are shown in Table 1, we focus solely on prompting LLMs to identify relevant columns and applicable methods, evaluating the accuracy of their selections instead of the computational results, as LLMs’ abilities in precise calculations are not the central focus of this study. In addition, we evaluate performance of information extraction using Entity-Attributed F1 (EA-F1). This is an F1 score applied to the predicted vs. gold sets of the (entity, atrribution, value). All three elements in the tuple must exactly match the tuple in the ground truth to be marked correct.
4.2 Results and Analysis
@@
Main result. GPT-4 + RAG achieved superior accuracy (59.3%), outperforming the second-ranked model (FT Llama-3-Instruct: 55.6% ) by a statistically significant margin. Notably, GPT-4 + RAG excelled in relational (68.7%) and comparative (76.3%) queries, likely due to its superior contextual understanding. However, all models exhibited markedly lower accuracy in statistical queries (GPT-4 + RAG: 41.0%), suggesting inherent challenges in numerical reasoning. In our evaluation, we focused on analyzing the capability of LLMs to extract question-related data. This assessment aimed to understand how well these sophisticated models can organize and present data for the question. The result is shown in Table 4. These results underscore the critical role of information extraction architectures in mitigating hallucinations and grounding outputs in factual data. Introducing RAG significantly improves overall performance, particularly in comparison tasks, while fine-tuning LLaMA-3-Instruct alone does not yield substantial gains without RAG. On MEBench, open-source models like LLaMA-3-Instruct, even with RAG, can’t match proprietary models like GPT-4, which achieves a 59.3% accuracy compared to LLaMA-3-Instruct’s 32.5%.
Fine-grained Performance on Sub-tasks. Figure 3 shows that vanilla LLMs perform well in correlation analysis and descriptive relationship tasks, while RAG significantly improves intercomparison and superlative tasks. However, neither fine-tuning nor RAG overcomes challenges in variance analysis and aggregation tasks, while GPT-4 + RAG achieves accuracy of 15.3% and 32.1%.
Entity density Analysis. As we can see from Table 3, our experiments underscore the impact of entity density on model performance in MEQA tasks. This phenomenon arises because higher entity densities amplify two critical challenges inherent to MEQA systems: (1) Semantic ambiguity due to overlapping relational predicates among entities (e.g., distinguishing "Paris [person]" vs. "Paris [location]" within narrow contexts), and (2) computational overhead in attention-based architectures attempting parallel reasoning over entangled entity-attribution pairs (e.g. transformer self-attention weights saturate under dense cross-entity dependencies).
  • Low Entity Density: Models generally performed well in low-density scenarios. The simplicity of context allowed for accurate entity recognition and minimal ambiguity.
@@
5 Limitations
While MEBench provides a comprehensive framework for evaluating cross-document multi-entity reasoning, our work has several limitations that warrant further investigation. Although MEBench covers eight distinct reasoning types across three broad categories, real-world MEQA scenarios may involve even more intricate combinations of logical, temporal, or causal dependencies. The current benchmark does not explicitly model dynamic or time-sensitive entity interactions, which could limit its applicability to domains like financial forecasting or event-driven narratives. The benchmark relies on a curated collection of documents to ensure controlled evaluation. While this design choice minimizes noise, it may not fully replicate the challenges of real-world environments where documents vary widely in quality, redundancy, and structure. Future iterations could incorporate noisy or incomplete data sources to better simulate practical scenarios. While the Entity-Attributed F1 (EA-F1) metric rigorously assesses entity-level correctness and attribution validity, it prioritizes factual precision over semantic coherence. This may undervalue partially correct answers that demonstrate valid reasoning chains but contain minor factual inaccuracies. A hybrid evaluation framework combining EA-F1 with human judgment could provide a more holistic assessment.
6 Conclusion
In this study, we have comprehensively addressed the significant challenges that Multi-entity Question Answering (MEQA) poses to LLMs and RAG systems. The limitations of existing methods in handling cross-document aggregation, especially when dealing with entity-dense questions, have been clearly identified and analyzed. We introduced MEBench, a groundbreaking multi-document, multi-entity benchmark. Our experiments on state-of-the-art LLMs such as GPT-4 and Llama-3, along with RAG pipelines, have shed light on the critical limitations of these advanced models. The fact that even these leading models achieve only 59% accuracy on MEBench underscores the magnitude of the challenges in MEQA. MEBench has effectively highlighted the systemic weaknesses in current LLM frameworks. These weaknesses serve as valuable insights for future research directions. For instance, the need for improved algorithms to retrieve and consolidate fragmented information from heterogeneous sources is evident. Additionally, there is a pressing need to develop more robust entity-aware QA architectures that can better handle the complexities of MEQA.
7 Acknowledgment
This work is supported by Guangdong provincial project 2023CX10X008.
@@ Appendix A Appendix
Table 5: Example Topics and Their Entities Attributions.
@@
Table 6: Template example for questions generated by the LLM (GPT-4).
@@
A.1 Methodology for composite SPARQL Generation via Iterative Semantic Refinement
A.1.1 Initial Query Parsing Using GPT-4
Entity Misalignment: Incorrect Wikidata Q-ID assignments due to lexical ambiguity (e.g., "Java" as programming language vs. Indonesian island)
A.1.2 Semantic Validation Layer
(a) Structured Knowledge Anchoring
The system interfaces with the Wikipedia API through programmatic endpoints that map surface forms to canonical entities.
(b) Neural-Semantic Disambiguation Module
GPT-4 serves as our semantic analysis engine, performing three key operations:
  • Contextual disambiguation using entity linking algorithms enhanced by Wikifier-style mention detection.
  • Property type validation against Wikidata’s ontology constraints (rdf:type, owl:ObjectProperty).
  • Temporal scope verification for time-sensitive queries requiring qualifiers like P585 [point in time].
A.1.3 Iterative Refinement Protocol
  • Executing candidate SPARQL on the Wikidata Query Service endpoint.
  • Analyzing result cardinality and type consistency.
  • Applying constraint satisfaction heuristics:
FILTER (?population > 1e6 && ?country IN wd:Q30) # Example numerical/entity constraints
Each iteration tightens precision metrics until meeting termination criteria defined by either:
$\frac{|ValidResults_{t}|}{|TotalResults_{t}|}\geq\theta_{precision}\quad$
$(\theta=0.98\;\text{empirically})$
or maximum iteration thresholds.
A.1.4 Final Query Synthesis
A.2 Optimization
Model Selection.
Model selection is straightforward yet highly effective for optimization Liu et al. (2024). Our system comprises multiple tasks, necessitating the selection of the most suitable model for different tasks. For basic tasks, more affordable and faster LLMs can suffice, while utilization of the most advanced LLMs is essential in more complex tasks to ensure optimal performance. Specifically, our system employs powerful yet resource-intensive GPT-4 for tasks such as semantic analysis or generation of table schemas and SQL queries. In contrast, for more basic information extraction, we utilize open-source Mistral-7B, thereby achieving a balance between cost efficiency and functional performance.
LLM Input/Output Control
SplitWise Patel et al. (2023) shows that LLM inference time is generally proportional to the size of input and output tokens. Since GPT models decide the cost based on the input token, we try to minimize the input of large models. Meanwhile, we use the instructive prompt to reduce the size of the outputs generated by LLM without changing the quality of these outputs. The example of prompt is in Appendix A.2.1.
A.2.1 Prompt for Output Control
A.2.2 Quality Control
Question Templates.
The use of templates ensures that every question is crafted with a clear structure, making it easier for respondents to understand and answer them accurately. For relationship and complex statistic questions we turn the questions in a closed-ended style, as they require a specific response of either "yes" or "no", which make the answer in a standardized format. The examples of Question Templates is in the Appendix 6.
Question Refinement.
Manual review.
We assess the questions for accuracy, ensuring they are factually correct and relevant to our purpose. Manual reviews can also provide insights into whether the questions are likely to effectively elicit the intended information from answers, thereby contributing to the reliability and validity of the benchmark.
A.3 Tables